# X-Boundary: Establishing Exact Safety Boundary to Shield LLMs from Multi-Turn Jailbreaks without Compromising Usability

### v2

@@ B.4 Evaluations
  • ActorAttack (Ren et al., 2024b): Emphasizes role-playing scenarios to gradually induce harmful behavior. The multi-turn queries in SafeMTData_Attack_600 (Ren et al., 2024b) are used to attack victim models, and HarmBench classifier (Mazeika et al., 2024) is used to judge whether the attack is successful.
  • RedQueen (Jiang et al., 2024): Focuses on dynamic prompt engineering with iterative refinements. We use the template of RedQueen to generate 600 test data based on HarmBench, and use HarmBench classifier as the judge model.
  • Crescendo (Russinovich et al., 2024): Includes gradually escalating attacks that push the model to produce harmful content over multiple turns. GPT-3.5-turbo is used as the attack model and GPT-4o is utilized as the judge model.
@@
  • XSTest (Röttger et al., 2023): Examines model responses to boundary-case prompts involving sensitive but potentially valid information.
  • OKTest (Shi et al., 2024): Evaluates whether the model declines benign questions in real-world scenarios.
  • OR-Bench (Cui et al., 2024): Explicitly measures over-refusal rates on a suite of harmless queries.
  • PHTest: Comprises prompts that may look suspicious but are legitimately safe for the model to address.
@@
  • MMLU: A broad measure of knowledge in diverse domains.
  • GSM8K: A math reasoning benchmark to test step-by-step problem solving.
  • HumanEval (Chen et al., 2021): Assesses code generation capability, crucial for real-world AI applications.
@@
  • Over-Refusal Rate : The fraction of benign prompts that the model incorrectly refuses to answer. A lower over-refusal rate signifies better usability.
  • General Performance : We measure the model’s utility on standard benchmarks (MMLU, GSM8K, HumanEval) to ensure that defensive measures do not degrade essential capabilities. A higher score indicates stronger performance on domain knowledge, reasoning, or code generation.
Appendix C Theoretical Analysis of X-Boundary
@@
Proposition 1 in this paper is an application of Proposition 13 in (Weed & Bach, 2019).
Definition 1 ((Weed & Bach, 2019)).
@@
Proposition 2 (Proven in (Weed & Bach, 2019)).
@@
$\mathrm{Var}_{m}(\phi_{\#}\mu)=\mathbb{E}_{S,\tilde{S}\sim\mu^{m}}[{\mathcal{W}}_{1}(\phi_{\#}\mu_{S},\phi_{\#}\mu_{\tilde{S}})]\leq 2\mathbb{E}_{S\sim\mu^{m}}[{\mathcal{W}}_{p}^{p}(\mu,\mu_{S})]\leq 24\sqrt{\frac{n}{m}}.$ (3)
@@
Appendix D Case Study
@@
Figure 17: Multi-Turn Attack example from Qwen2.5-7B-Chat on Crescendo.

### v3

@@ D.5 Evaluations
  • ActorAttack actor_attack: Emphasizes role-playing scenarios to gradually induce harmful behavior. The multi-turn queries in SafeMTData_Attack_600 actor_attack are used to attack victim models, and HarmBench classifier harmbench is used to judge whether the attack is successful.
  • RedQueen red_queen: Focuses on dynamic prompt engineering with iterative refinements. We use the template of RedQueen to generate 600 test data based on HarmBench, and use HarmBench classifier as the judge model.
  • Crescendo crescendo: Includes gradually escalating attacks that push the model to produce harmful content over multiple turns. GPT-3.5-turbo is used as the attack model and GPT-4o is utilized as the judge model.
@@
  • XSTest xstest: Examines model responses to boundary-case prompts involving sensitive but potentially valid information.
  • OKTest oktest: Evaluates whether the model declines benign questions in real-world scenarios.
  • OR-Bench orbench: Explicitly measures over-refusal rates on a suite of harmless queries.
  • PHTest phtest: Comprises prompts that may look suspicious but are legitimately safe for the model to address.
@@
  • MMLU mmlu: A broad measure of knowledge in diverse domains.
  • GSM8K gsm8k: A math reasoning benchmark to test step-by-step problem solving.
  • HumanEval human_eval: Assesses code generation capability, crucial for real-world AI applications.
@@
  • Over-Refusal Rate (ORR): The fraction of benign prompts that the model incorrectly refuses to answer. A lower over-refusal rate signifies better usability.
  • General Capability: We measure the model’s utility on standard benchmarks (MMLU, GSM8K, HumanEval) to ensure that defensive measures do not degrade essential capabilities. A higher score indicates stronger performance on domain knowledge, reasoning, or code generation.
Appendix E Theoretical Analysis of X-Boundary
@@
Proposition 1 in this paper is an application of Proposition 13 in weed2017sharp.
Definition 2 (weed2017sharp).
@@
Proposition 3 (Proven in weed2017sharp).
@@
[…]
$\leq 2\mathbb{E}_{S\sim\mu^{m}}[{\mathcal{W}}_{p}^{p}(\mu,\mu_{S})]\leq 24\sqrt{\frac{n}{m}}.$ (11)
@@
Appendix F Case Study
@@
Figure 20: Multi-Turn Attack example from Qwen2.5-7B-Instruct on Crescendo.